# Revisiting Pre-trained Language Models for Vulnerability Detection

### v1

@@ III-G2 Hyperparameters
RobertaClassificationHead [64] is used as classifier when fine-tuning SLMs, which comprises a feedforward neural network with two linear layers (hidden_size=768), dropout regularization (rate=0.1), and a tanh activation function. The AdamW optimizer [65] is employed along with a linear learning rate schedule with a warm-up ratio set to 10% of total training steps. When fine-tuning LLMs, we use AutoModelForSequenceClassification to load base model and add LORA adapters to all linear layers of base model. We configure LoRA adapter with a rank of 64, an alpha of 16, and 0.05 dropout rate.For all experiments, we use a batch size of 32, 10 epochs, and a learning rate of 2e-5.
When prompting LLMs, we set the top-p value to 0.9, the temperature to 0, and limit the maximum number of new tokens to 10. These settings are chosen to ensure that the model produces concise and deterministic responses.
IV Research Questions (RQs) and Findings
@@
TABLE IV: Performance comparison of Code SLMs and LLMs on the self-collected evaluation dataset
@@
Fine-tuning-based evaluation
@@
42.71
76.06
51.71
52.52
59.58
62.77
5.71
8.83
7.85
2.99
4.93
4.14
62.89
42.14
74.47
42.14
77.02
51.06
IV-A RQ1: How do PLMs perform across various training and testing configurations?
Due to inconsistent experimental settings in existing VD research, it is difficult to compare evaluation results across studies, often leading to conflicting conclusions. For example, some studies fine-tune PLMs on imbalanced training sets and evaluate them on imbalanced test sets [33, 1], while others fine-tune on balanced training sets and evaluate on either imbalanced [10] or balanced test sets [13]. To investigate the impact of different experimental settings on model training, we fine-tune SLMs under both imbalanced and balanced training conditions. We then evaluate their generalizability using two imbalanced test sets with different distributions. We deliberately avoid using balanced test settings, as real-world security datasets are inherently imbalanced. Evaluating a model, regardless of its training setup, solely on balanced data risks significantly overestimating its precision.
From Table III, we observe that under imbalanced training settings, the average TNR of SLMs on both test sets exceeds 99%, whereas the average recall remains below 10%. This discrepancy is attributed to the overwhelming number of non-vulnerable functions compared to vulnerable ones, which biases the fine-tuned SLMs toward predicting negatives. Under balanced training settings, the average TNR of SLMs on the reconstructed test set is 73.5%, the average recall is 75.2%, and the average balanced accuracy is 74.4%. These suggest that SLMs fine-tuned under balanced settings are able to fairly learn the representations of both vulnerable and non-vulnerable functions, thereby achieving a better trade-off between recall and TNR. In particular, we observe that the F1 score tends to favor models trained on highly imbalanced data, creating a misleading impression that these models perform better under imbalanced training settings. In contrast, balanced accuracy is the average of recall and TNR, so any bias toward either class will equally affect the balanced accuracy.
Among SLMs fine-tuned under balanced training settings, PDBERT performs the best with its balanced accuracy averaging about 3.2% to 4.7% higher than that of the others on both test sets. UniXCoder and GraphCodeBERT follow behind, while CodeBERT exhibits the lowest performance. PDBERT’s superior performance is attributed to its incorporation of both control and data dependency prediction tasks during pre-training, which enables the model to capture vulnerability-related patterns by analyzing dependencies within the code, thereby enhancing its detection ability. In contrast, UniXCoder and GraphCodeBERT, which are exposed to only single aspects of data flow or AST during pre-training, show somewhat lower performance. Without learning code structure during pre-training, CodeBERT presents limited ability to accurately distinguish between vulnerable and non-vulnerable functions.
[…]
IV-B RQ2: How do PLMs of varying parameter scales perform under both fine-tuning and prompt engineering?
Given the strong performance of LLMs in NLP and software engineering, and the efficiency of prompt engineering compared to fine-tuning, most prior studies prefer prompting-based evaluations for VD, leaving comprehensive evaluations of fine-tuned LLMs with parameter sizes up to 34B underexplored. Additionally, Code SLMs, pre-trained on structure-aware code tasks, are expected to be better suited for VD due to their ability to capture complex code structures and dependencies. With significantly fewer parameters than LLMs, Code SLMs can be fully fine-tuned on VD datasets with minimal computational cost. To this end, we compare Code SLMs and LLMs on VD tasks under both fine-tuning and prompt settings, analyzing the trade-off between performance and efficiency.
Specifically, in prompting-based evaluations, for each LLM, we apply four types of prompts (Section III-E) under both zero-shot and few-shot settings, and report the best testing results. The fine-tuning-based evaluations are conducted in balanced reconstruction training set (best setting shown in Section IV-A). For Code SLMs, we perform full fine-tuning. For LLMs, we apply parameter-efficient fine-tuning (PEFT) using LoRA [66], which has shown performance comparable to full fine-tuning. To avoid data leakage and ensure accurate evaluation, all PLMs are tested on our new self-collected dataset containing commits from October 2023, after the pretraining cutoff of all evaluated models. Full results and analysis are provided in Appendix A.
As shown in Table IV, the overall performance of fine-tuned PLMs significantly outperform that of prompting-based PLMs by about 15%. This suggests that, although few-shot examples or code structure and dependency information are provided during prompting to guide LLMs in understanding code, the complexity and diversity of vulnerability patterns still pose substantial challenges. In contrast, fine-tuning PLMs involves feeding the models a large amount of vulnerability-related code and guiding them through supervised classification to distinguish between vulnerable and non-vulnerable code, thereby enhancing their capabilities for VD.
In the prompting-based evaluation, GPT-4o Mini and DeepSeek-Coder-34B perform the best, confirming their strong performance in code-related tasks. As the parameter scale increases, the performance of LLMs also improves, validating that more code data fed during the pre-training phase enhances the models’ general understanding of code. This in turn aids in downstream VD tasks by allowing models to capture code logic and dependency relationships for detecting potential vulnerabilities. Notably, for CodeLlama and some relatively smaller LLMs (e.g., WizardCoder-15B and Starchat-beta-16B), their TNR is below 10%. As parameter scale increases, this issue is partially mitigated in WizardCoder, with its 34B model showing a 50% improvement in TNR compared to its 15B counterpart. However, for CodeLlama, despite maintaining a high recall, its TNR remains low. This may be due to the model’s oversensitivity to certain keywords in the prompts or code, leading to unreliability in VD task [7].
[…]
IV-C RQ3: How do PLMs perform across various types of vulnerabilities?
Figure 1: PLMs’ recall by vulnerability type
(ordered by decreasing frequency from left to right)
In this section, we evaluate the ability of PLMs to detect individual CWEs in our reconstructed dataset. We select representative models that perform well in Table IV from different parameter scales. As shown in Fig. 1, we observe that all PLMs achieve over 70% recall on average and perform particularly well on specific CWEs, such as CWE-119 (Improper Restriction of Operations within the Bounds of a Memory Buffer), CWE-399 (Resource Management Errors), CWE-189 (Numeric Errors), and CWE-787 (Out-of-bounds Write). Detecting these types of vulnerabilities often involves recognizing the absence of sanity checks within the code context. For example, CWE-119 typically arises when a program fails to validate boundary sizes. PLMs can relatively easily identify that by checking whether memory buffer accesses are properly guarded, such as through the use of if statements that ensure array indices remain within valid bounds.
However, it can be challenging for most PLMs to detect certain types of vulnerabilities, such as CWE-264 (Permissions, Privileges, and Access Controls) and CWE-476 (Null Pointer Dereference). Detecting these issues often requires complex dependency analysis or additional external information. For instance, identifying CWE-264 vulnerabilities may necessitate an understanding of the software’s permission control scheme to evaluate whether the permissions used in the code are appropriate. Information available within a single function is often insufficient for making an accurate prediction.
[…]
Answer-3: Leveraging the fundamental capabilities of code understanding acquired from pre-training on large-scale code corpora, PLMs generally perform well on vulnerabilities with clear structural patterns. However, they struggle with vulnerabilities that involve complex data and control dependencies, or those that require external information to be accurately identified. Notably, for certain complex vulnerabilities, PDBERT outperforms other models, suggesting that incorporating program dependencies during pre-training can better guide models in understanding code semantics. This suggests that future research could explore designing pre-training objectives tailored to specific code structures or vulnerability characteristics to further improve PLM performance in VD.
[…]
IV-D RQ4: How do PLMs perform under different code normalization rules?
Figure 2: PLMs’ performance under different code normalization rules (* marks model’s default normalization)
@@ IV-D RQ4: How do PLMs perform under different code normalization rules?
[…]
IV-E RQ5: Are PLMs capable of understanding abstracted vulnerability code?
Figure 3: PLMs’ performance before and after abstraction
For privacy concerns, developers may rename identifiers in their code when using PLMs for code analysis tasks. To examine whether PLMs can still understand vulnerable code without relying on textual information, we abstracted all identifiers and strings in the samples within the reconstructed test set. Specifically, we used Joern [68] to identify all variables, parameters, and strings within the functions, and replaced them with abstracted forms such as VAR0, PARAM0, and STRING0, where the trailing numbers distinguish different entities. We then fed both the original and abstracted functions into PLMs and compared their prediction results.
@@ IV-E RQ5: Are PLMs capable of understanding abstracted vulnerability code?
IV-F RQ6: Are PLMs capable of detecting vulnerable code after semantic-preserving transformation?
@@
Figure 4: PLMs’ performance before and after semantic-preserving transformations
@@
Effective PLMs should demonstrate robustness against semantic-preserving transformations for two key reasons. First, the widespread adoption of OSS has amplified the risk of software vulnerabilities, as insecure code is often cloned and reused with minor syntactic modifications during copy-paste by less experienced developers [69, 70]. Second, attackers may launch adversarial example attacks by deliberately applying SPTs to modify vulnerable code in ways that evade detection[71, 72]. However, detecting such vulnerabilities is challenging: although the underlying semantic flaws persist, the code syntactic patterns are altered.
@@
To evaluate whether PLMs can produce consistent predictions for vulnerable code clone/reuse, we apply four types of semantic-preserving transformations (i.e., conditional branch negation, conditional branch expansion, loop branch conversion, and relational operator reversal) to all functions in the reconstructed test set. Specifically, if a transformation rule matches $M$ positions in a function, we apply the rule to each position individually, generating $M$ transformed functions. In total, based on 23,144 original functions, we generate 48,182 semantically equivalent functions. We evaluate PLMs on both the original and transformed test sets and report their performance for each type of transformation accordingly.
@@
As shown in Fig. 4, the performance of PLMs degrades to varying degrees on different types of transformed data, whereas UniXCoder demonstrates strong robustness to code transformations with an average performance drop of only 0.1% after transformation. A reasonable explanation is that its pre-training phase incorporates a multi-modal contrastive learning task, which helps it maintain consistent predictions for perturbed positive samples. In particular, PLMs show the highest robustness to the conditional branch expansion transformation, with an average performance decrease of only 1.3% after transformation. The reason may be that breaking down compound conditions into separate branches provides more detailed logical steps while preserving the original semantic structure. However, PLMs exhibit the lowest robustness to the conditional branch negation transformation, with an average performance decrease of 15.8%. This may be because this transformation significantly changes the positions of code snippets, which could cause easily trackable data flows to be separated by unrelated code, making it difficult for PLMs to reliably extract long-distance dependencies. Additionally, PLMs have moderate robustness to relational operator reversal and loop branch conversion transformations. While reversed comparisons do not change the code logic, they might still pose challenges in model parsing due to variations in tokenization or the interpretation of relational operators. Similarly, although the syntactic differences between while and for loops are relatively small, initializing, condition-checking, and incrementing within a single line (as in for loops) versus being spread out (as in while loops) may introduce complexity in semantic parsing.
[…]
@@
IV-G RQ7: To what extent do implicit labeling errors caused by token truncation affect the performance of Code SLMs?
Figure 5: Code SLMs’ performance before and after slicing
@@
In addition to labeling errors during data collection, we identify an implicit labeling error that can result in unintentional data poisoning during token truncation. Due to the limited context window size (e.g., 512 tokens) of most Code SLMs, input functions are often truncated when fed into the models. This truncation prevents the models from accessing complete code context, forcing them to make predictions based on incomplete or less vulnerability-relevant code segments. Consequently, model training can be misled by these noisy labels, potentially compromising the model’s learning process.
@@
For example, in a memory leakage vulnerability-related commit (ID 0ddcff4 in Linux Kernel [73]), we find that due to truncation, the visible input portions of the vulnerable and non-vulnerable function appear identical to the Code SLMs. However, their ground-truth labels are opposite, meaning that the same inputs correspond to opposite labels. This will confuse model training, resulting in unsatisfactory model performance. By analyzing 5,480 patch pairs in PrimeVul, we find that 1,473 pairs (27%) have this issue. Although one naive way to avoid the labeling error caused by truncation is to simply exclude functions exceeding the maximum input size of Code SLMs, this limits the size of training data, leading to unsatisfactory model performance and making the learned models hard to deploy in real-world systems.
@@
Since not all statements in vulnerable functions are related to vulnerabilities, we perform code slicing to reduce the input length while preserving the core parts related to vulnerabilities as much as possible, thus minimizing the occurrence of labeling errors and enabling the model to learn more complete vulnerability patterns in functions. Specifically, we use tree-climber [74] to extracting both data and control flow of each patch pair in PrimeVul. To identify the core lines related to vulnerabilities, inspired by SySeVR [51], we begin with designating lines that include API function calls, array usage, pointer usage, and arithmetic expressions as anchor lines. Then, to ensure all relevant dependencies are extracted, we perform slicing bidirectionally by forwarding and backtracking from the anchored lines. Finally, to prevent the total token count for sliced code lines from exceeding the maximum input size of the Code SLMs (i.e., 512), we implement a length checker that stops slicing when the token limit is reached.
@@
After slicing, we fine-tune the Code SLMs on both original paired training set of PrimeVul and sliced training set, and evaluate the performance of Code SLMs on their corresponding paired test set. As shown in Fig. 5, Code SLMs fine-tuned on the sliced training set achieve an average performance improvement of 4% compared to those trained on the original paired training set. This supports our earlier hypothesis that token truncation introduces a substantial amount of poisoned data during training, ultimately degrading model performance. It is important to note that although state-of-the-art LLMs have a larger context window size to support function-level VD, the issue of labeling errors caused by token truncation still remains in cross-function or cross-file VD tasks.
[…]

### v3

@@ 3.7.2. Hyperparameters
RobertaClassificationHead (hf) is used as classifier when fine-tuning SLMs, which comprises a feedforward neural network with two linear layers (hidden_size=768), dropout regularization (rate=0.1), and a tanh activation function. The AdamW optimizer (adamw) is employed along with a linear learning rate schedule with a warm-up ratio set to 10% of total training steps. When fine-tuning LLMs, we use AutoModelForSequenceClassification to load base model and add LORA adapters to all linear layers of base model. We configure LoRA adapter with a rank of 64, an alpha of 16, and 0.05 dropout rate. For all experiments, we use a batch size of 32, 10 epochs, and a learning rate of 2e-5.
When prompting LLMs, we set the top-p value to 0.9, the temperature to 0. For GPT-5 Mini, we keep the temperature to default 1.0 as it cannot be modified. The reasoning effort is set to medium, and the verbosity is low because we only focus on the correctness of the binary label prediction. These settings are chosen to ensure that the model produces concise and deterministic answers.
4. Research Questions (RQs) and Findings
4.1. RQ1: How do PLMs perform across various training and testing configurations?
Due to inconsistent experimental settings in existing VD research, it is difficult to compare evaluation results across studies, often leading to conflicting conclusions. For example, some studies fine-tune PLMs on imbalanced training sets and evaluate them on imbalanced test sets (primevul; diversevul), while others fine-tune on balanced training sets and evaluate on either imbalanced (yinxin) or balanced test sets (davidlo). To investigate the impact of different experimental settings on model training and testing, we fine-tune Code SLMs under both imbalanced and balanced training conditions (LLMs are excluded because they share the same conclusion as those of Code SLMs). We then evaluate their generalizability using two imbalanced test sets with different distributions. We deliberately avoid using balanced test settings, as real-world security data is inherently imbalanced. Evaluating a model, regardless of its training setup, solely on balanced data risks significantly overestimating its precision.
From Table 4, we observe that under imbalanced training settings, the average TNR of Code SLMs on both test sets exceeds 99%, whereas the average recall remains below 10%. This discrepancy is attributed to the overwhelming number of non-vulnerable functions compared to vulnerable ones, which biases the fine-tuned Code SLMs toward predicting negatives. Under balanced training settings, the average TNR of Code SLMs on the reconstructed test set is 73.5%, the average recall is 75.3%, and the average balanced accuracy is 74.4%. These suggest that Code SLMs fine-tuned under balanced settings are able to fairly learn the representations of both vulnerable and non-vulnerable functions, thereby achieving a better trade-off between recall and TNR. In particular, we observe that the F1 score tends to favor models trained on highly imbalanced data, creating a misleading impression that these models perform better under imbalanced training settings. In addition, all PLMs show a low precision, which is because the imbalanced distribution of real-world vulnerability data (non-vulnerable data is far more than vulnerable data) results in the number of FPs being much larger than the number of TPs. This also explains why we chose balanced accuracy as our main metric; it considers both recall and TNR, so any bias toward either class will equally affect the result.
Among Code SLMs fine-tuned under balanced training settings, PDBERT performs the best with its balanced accuracy averaging about 3.2% to 4.7% higher than that of the others on both test sets. UniXCoder and GraphCodeBERT follow behind, while CodeBERT exhibits the lowest performance. PDBERT’s superior performance is attributed to its incorporation of both control and data dependency prediction tasks during pre-training, which enables the model to capture vulnerability-related patterns by analyzing dependencies within the code, thereby enhancing its detection ability. In contrast, UniXCoder and GraphCodeBERT, which are exposed to only single aspects of data flow or AST during pre-training, show somewhat lower performance. Without learning code structure during pre-training, CodeBERT presents limited ability to accurately distinguish between vulnerable and non-vulnerable functions.
Answer-1: (1) Selecting appropriate experimental settings is essential for accurately evaluating the capabilities of PLMs for VD. Fine-tuning PLMs under balanced training settings helps mitigate prediction bias. Evaluating PLMs on imbalanced test sets ensures a realistic performance assessment in practical scenarios. (2) Among existing Code SLMs, PDBERT achieves the best performance, demonstrating that guiding models to learn code dependency relationships during pre-training effectively enhances VD. (3) Evaluating fine-tuned PLMs on test data sourced from the same distribution as the training data leads to overestimated results that fail to reflect the models’ real ability for VD in real-world scenarios.
@@ 4.1. RQ1: How do PLMs perform across various training and testing configurations?
Table 5. Performance comparison of Code SLMs and LLMs on the self-collected evaluation dataset
@@
Fine Tuning-based Evaluation
@@
52.17
61.11
51.92
@@
5.29
7.11
7.22
2.78
3.78
3.80
52.79
58.82
73.99
52.16
65.17
51.35
4.2. RQ2: How do PLMs of varying parameter scales perform under both fine-tuning and prompt engineering?
Given the strong performance of LLMs in NLP and software engineering, and the efficiency of prompt engineering compared to fine-tuning, most prior studies prefer prompting-based evaluations for VD, leaving comprehensive evaluations of fine-tuned LLMs with parameter sizes up to 34B underexplored. Additionally, Code SLMs, pre-trained on structure-aware code tasks, are expected to be better suited for VD due to their ability to capture complex code structures and dependencies. With significantly fewer parameters than LLMs, Code SLMs can be fully fine-tuned on VD datasets with less computational cost. To this end, we compare Code SLMs and LLMs on VD tasks under both fine-tuning and prompt settings, analyzing the trade-off between performance and efficiency.
Specifically, in prompting-based evaluations, for each LLM, we apply four types of prompts (Section 3.5) under both zero-shot and few-shot settings, and report the best testing results. The fine-tuning-based evaluations are conducted in balanced reconstruction training set (best setting shown in Section 4.1). For Code SLMs, we perform full fine-tuning. For LLMs, we apply parameter-efficient fine-tuning (PEFT) using LoRA (hu2022lora), which has shown performance comparable to full fine-tuning. To avoid data leakage and ensure accurate evaluation, all PLMs are tested on our new self-collected dataset containing commits from October 2023, after the pretraining cutoff of all evaluated models.
As shown in Table 5, the overall performance of fine-tuned PLMs significantly outperform that of prompting-based PLMs by about 14%. This suggests that, although few-shot examples or code structure and dependency information are provided during prompting to guide LLMs in understanding code, the complexity and diversity of vulnerability patterns still pose substantial challenges. In contrast, fine-tuning PLMs involves feeding the models a large amount of vulnerability-related code and guiding them through supervised classification to distinguish between vulnerable and non-vulnerable code, thereby enhancing their capabilities for VD.
In the prompting-based evaluation, GPT-4o Mini and DeepSeek-Coder-34B perform the best, confirming their strong performance in code-related tasks. As the parameter scale increases, the performance of LLMs also improves, validating that more code data fed during the pre-training phase enhances the models’ general understanding of code. This in turn aids in downstream VD tasks by allowing models to capture code logic and dependency relationships for detecting potential vulnerabilities. Notably, for CodeLlama and some relatively smaller LLMs (e.g., WizardCoder-15B and Starchat-$\beta$-16B), their TNR is below 10%. As parameter scale increases, this issue is partially mitigated in WizardCoder, with its 34B model showing a 52% improvement in TNR compared to its 15B counterpart. However, for CodeLlama, despite maintaining a high recall, its TNR remains low. This may be due to the model’s oversensitivity to certain keywords in prompts, leading to unreliability in VD (benjamin2023).
Figure 2. Performance comparison of PLMs on the self-collected dataset beyond the GPT-5 Mini cutoff date:
Performance Evaluation of GPT-5 Mini. Recent advancements in PLMs with reasoning capabilities have shown strong performance in the field of software engineering. To complement this and highlight the timeliness of our findings, we include the SOTA reasoning model GPT-5 Mini, which has demonstrated leading performance in code benchmarks such as SWE-Bench (swebench).
As the knowledge cutoff date for GPT-5 Mini is June 2024, we select a subset of the self-collected evaluation data with commit dates starting from July 2024 to ensure accuracy. For a fair comparison, we also report the best balanced accuracy of other PLMs on the same filtered subset. As shown in Fig. 2, while GPT-5 Mini shows a relative improvement in performance compared to GPT-4o Mini, it still fails to outperform most fine-tuned PLMs. This once again underscores the importance of fine-tuning for improving PLMs’ capabilities for VD.
Answer-2: (1) Existing research tends to favor applying prompting-based PLMs for VD, while overlooking a comprehensive evaluation of fine-tuned PLMs across parameter scales. While prompt engineering is efficient, its performance on VD tasks falls far short of that achieved through fine-tuning. This highlights the importance of balancing performance and efficiency in practical applications. (2) Scaling law generally holds in both fine-tuning and prompting-based evaluations, but its impact varies across PLMs. (3) Compared to pre-training solely on large code corpora with standard language modeling, incorporating code-specific optimization during pre-training proves more beneficial for achieving strong performance on VD tasks, which demand a deep understanding of code logic and dependencies.
4.3. RQ3: How do PLMs perform across various types of vulnerabilities?
Figure 3. PLMs’ recall by vulnerability type
(ordered by decreasing frequency from left to right)
In this section, we evaluate the ability of PLMs to detect individual CWEs on our reconstructed dataset. To ensure the representativeness of conclusions, in Sections 4.3-4.6, we select representative models that perform well in Table 5 from different parameter scales. As shown in Fig. 3, we observe that all PLMs achieve over 70% recall on average and perform particularly well on specific CWEs, such as CWE-119 (Improper Restriction of Operations within the Bounds of a Memory Buffer), CWE-399 (Resource Management Errors), CWE-189 (Numeric Errors), and CWE-787 (Out-of-bounds Write). Detecting these types of vulnerabilities often involves recognizing the absence of sanity checks within the code context. For example, CWE-119 typically arises when a program fails to validate boundary sizes. PLMs can relatively easily identify that by checking whether memory buffer are properly accessed (e.g., adding boundary checks before sensitive operations).
However, it can be challenging for most PLMs to detect certain types of vulnerabilities, such as CWE-264 (Permissions, Privileges, and Access Controls) and CWE-476 (Null Pointer Dereference). Detecting these issues often requires complex dependency analysis or additional external information. For instance, identifying CWE-264 vulnerabilities may necessitate an understanding of the software’s permission control scheme to evaluate whether the permissions used in the code are appropriate. A null pointer dereference vulnerability also requires cross-procedure information, which is not available within a single function, to make an accurate prediction.
Answer-3: Leveraging the fundamental capabilities of code understanding acquired from pre-training on large-scale code corpora, PLMs generally perform well on vulnerabilities with obvious vulnerability patterns. However, they struggle with vulnerabilities that involve complex data and control dependencies, or those that require contextual information to be accurately identified. Notably, for certain complex vulnerabilities, PDBERT outperforms other models, suggesting that incorporating program dependencies during pre-training can better guide models in understanding code semantics. This suggests that future research could explore designing pre-training objectives tailored to specific code structures or vulnerability characteristics to further improve PLM performance in VD.
4.4. RQ4: How do PLMs perform under different code normalization rules?
Figure 4. PLMs’ performance under different code normalization rules (* marks model’s default normalization):
@@ 4.4. RQ4: How do PLMs perform under different code normalization rules?
Answer-4: Inconsistency in code normalization during training and testing leads to significant performance degradation of PLMs, indicating that they lack robustness to minor perturbations in code format.
4.5. RQ5: Are PLMs capable of understanding abstracted vulnerability code?
Figure 5. PLMs’ performance before and after abstraction:
For privacy concerns, developers may mask the name of identifiers in their code when using PLMs for code analysis tasks. To examine whether PLMs can still understand vulnerable code without relying on textual information, we abstract all identifiers and strings in the code within the reconstructed test set. Specifically, we use Joern (Joern) to identify all variables, parameters, and strings within the functions, and replace them with abstracted forms such as VAR0, PARAM0, and STRING0, where the trailing numbers distinguish different entities. We then input both the original and abstracted functions into PLMs and compared their prediction results.
Answer-5: PLMs are to some extent robust to code abstraction, which suggests that their vulnerability predictions do not rely solely on textual content. Instead, the fundamental code understanding capabilities learned from large code corpora, or the ability to analyze program dependency relationships acquired during pre-training, help them comprehend code logic and structural information for VD.
@@ 4.5. RQ5: Are PLMs capable of understanding abstracted vulnerability code?
4.6. RQ6: Are PLMs capable of detecting vulnerable code after semantic-preserving transformation?
@@
Figure 6. PLMs’ performance before and after semantic-preserving transformations:
@@
Effective PLMs should demonstrate robustness against semantic-preserving transformations for two key reasons. First, the widespread adoption of OSS has amplified the risk of software vulnerabilities, as insecure code is often cloned and reused with minor syntactic modifications during copy-paste by less experienced developers (VGRAPH; cpminer). Second, attackers may launch adversarial example attacks by deliberately applying semantic-preserving transformations to modify vulnerable code in ways that evade detection (yang2022natural; zhang2023challenging). However, detecting such vulnerabilities is challenging: although the underlying semantic flaws persist, the code syntax are altered.
(a) Conditional branch negation
(b) Relational operator reversal
(c) Conditional branch expansion
(d) Loop branch conversion
Figure 7. Semantic-preserving transformations:
@@
To evaluate whether PLMs can produce consistent predictions for vulnerable code clone/reuse, as shown in Fig. 7, we apply four types of semantic-preserving transformations (i.e., conditional branch negation, conditional branch expansion, loop branch conversion, and relational operator reversal) to all functions in the reconstructed test set. Specifically, if a transformation rule matches $M$ positions in a function, we apply the rule to each position individually, generating $M$ transformed functions. In total, based on 23,144 original functions, we generate 48,182 semantically equivalent functions. We evaluate PLMs on both the original and transformed test sets and report their performance for each type of transformation.
@@
As shown in Fig. 6, the performance of PLMs degrades to varying degrees on different types of transformed data, whereas UniXCoder demonstrates strong robustness to code transformations with an average performance drop of only 0.1% after transformation. A reasonable explanation is that its pre-training phase incorporates a multi-modal contrastive learning task, which helps it maintain consistent predictions for perturbed positive samples. In particular, PLMs show the highest robustness to the conditional branch expansion transformation, with an average performance decrease of only 1.3% after transformation. A possible reason is that breaking down compound conditions into separate branches provides more detailed logical steps while preserving the original semantic structure. However, PLMs exhibit the lowest robustness to the conditional branch negation transformation, with an average performance decrease of 15.8%. This may be because this transformation significantly changes the positions of code snippets, which could cause easily trackable data flows to be separated by unrelated code, making it difficult for PLMs to reliably extract long-distance dependencies. Additionally, PLMs have moderate robustness to relational operator reversal and loop branch conversion transformations. While reversed comparisons do not change the code logic, they might still pose challenges in model parsing due to variations in tokenization or the interpretation of relational operators. Similarly, although the syntactic differences between while and for loops are relatively small, initializing, condition-checking, and incrementing within a single line (as in for loops) versus being spread out (as in while loops) may introduce complexity in semantic parsing.
Answer-6: Most PLMs exhibit varing degrees of performance drop in semantic-preserving code, indicating they are not yet reliable for detecting potential vulnerable code clones and reuse in real-world scenarios. In contrast, UniXCoder mitigates the impact of transformed samples and maintains consistent performance by introducing multi-modal contrastive learning during pre-training. This highlights the potential for future VD research to design pre-training tasks that enhance adversarial robustness, thereby improving resilience to code transformations.
@@
4.7. RQ7: To what extent do implicit labeling errors caused by token truncation affect the performance of Code SLMs?
Figure 8. Code SLMs’ performance before and after slicing:
@@
In addition to labeling errors during data collection, we identify an implicit labeling error that can result in unintentional data poisoning during token truncation. Due to the limited context window size (e.g., 512 tokens) of most Code SLMs, input functions are often truncated when fed into the models. This truncation prevents the models from accessing complete code context, forcing them to make predictions based on incomplete or less vulnerability-relevant code segments. Consequently, model training can be misled by these noisy labels, potentially compromising the training process.
Figure 9. Limited visible part caused by truncation:
@@
For example, Fig. 9 shows a memory leakage vulnerability-patched commit (ID 0ddcff4 in Linux Kernel (Linux_commit)). Due to truncation, the visible input portions of the vulnerable and non-vulnerable function appear identical to the Code SLMs. However, their ground-truth labels are opposite, meaning that the same inputs correspond to opposite labels. This will confuse model training, resulting in unsatisfactory model performance. By analyzing 5,480 patch pairs in PrimeVul, we find that 1,473 pairs (27%) have this issue. Although one naive way to avoid the labeling error caused by truncation is to simply exclude functions exceeding the maximum input size of Code SLMs, this limits the size of training data, leading to unsatisfactory model performance and making the learned models hard to deploy in real-world systems.
@@
Since not all statements in vulnerable functions are related to vulnerabilities, we perform code slicing to reduce the input length while preserving the core parts related to vulnerabilities as much as possible, thus minimizing the occurrence of labeling errors and enabling the model to learn more complete vulnerability patterns in functions. Specifically, we use tree-climber (treeclimber) to extracting both data and control flow of each patch pair in PrimeVul. To identify the core lines related to vulnerabilities, inspired by SySeVR (SySeVR), we begin with designating lines that include API function calls, array usage, pointer usage, and arithmetic expressions as anchor lines. Then, to ensure all relevant dependencies are extracted, we perform slicing bidirectionally by forwarding and backtracking from the anchored lines. Finally, to prevent the total token count for sliced code lines from exceeding the maximum input size of the Code SLMs (i.e., 512), we implement a length checker that stops slicing when the token limit is reached.
@@
After slicing, we fine-tune the Code SLMs on both original paired training set of PrimeVul and sliced training set, and evaluate the performance of Code SLMs on their corresponding paired test set. As shown in Fig. 8, Code SLMs fine-tuned on the sliced training set achieve an average performance improvement of 4% compared to those trained on the original paired training set. This supports our earlier hypothesis that token truncation introduces a substantial amount of poisoned data during training, ultimately degrading model performance. It is important to note that although SOTA LLMs have a larger context window size to support function-level VD, the issue of labeling errors caused by token truncation still remains in repo-level VD tasks, which requires an understanding of the entire codebase in real-world projects.